OpenStudy (anonymous):

What is the solution to the linear equation? d – 10 – 2d + 7 = 8 + d – 10 – 3d a. d = –5 b. d = –1 c. d = 1 d. d = 5

OpenStudy (anonymous):

@Night-Watcher

OpenStudy (anonymous):

@bibby

OpenStudy (anonymous):

Kik ?

OpenStudy (bibby):

combine the like terms

OpenStudy (anonymous):

I did

OpenStudy (bibby):

d – 10 – 2d + 7 = 8 + d – 10 – 3d so what do we have after we combine like terms?

OpenStudy (bibby):

god dammit

OpenStudy (bibby):

d-10-2d+7 = 8 + d - 10 - 3d -d-3=8+d-10-3d

OpenStudy (anonymous):

Whats next

OpenStudy (bibby):

do the same thing to the right side of the equation

OpenStudy (anonymous):

Its C